OpenStudy (anonymous):

What is the simplified form of the quantity x^2 -6 over the quantity x+4 ?

OpenStudy (anonymous):



OpenStudy (anonymous):

is it x² - 6 or x² - 16 ?

OpenStudy (anonymous):

16 sorry

OpenStudy (anonymous):

\[\frac{x^2 - 16}{x + 4} = \frac{\left( x + 4 \right)\left( x - 4 \right)}{x + 4} = x - 4\]

OpenStudy (anonymous):

the restriction would be positive 4?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

negative 4, because\[x + 4 \neq 0\]

OpenStudy (anonymous):

so the answer is x-4 with a restriction of -4

OpenStudy (anonymous):

but you don't need any restrictions for the final answer because the simplied form doesn't have any denominator